# Probing quantum properties of black holes with a Floquet-driven optical lattice simulator

### v1

@@ Abstract
In the curved spacetime of a black hole, quantum physics gives rise to distinctive effects such as Hawking radiation. Here, we present a scheme for an analogue quantum simulation of $(1+1)$-dimensional black holes using ultracold atoms in a locally Floquet-driven 1D optical lattice. We show how the effective dynamics of the driven system can generate a position-dependent tunnelling profile that encodes the curved geometry of the black hole. Moreover, we provide a simple and robust scheme to determine the Hawking temperature of the simulated black hole based solely on on-site atom population measurements. Finally, we show how this scheme can be directly applied to simulate $(2+1)$D black holes by utilising 2D optical lattices. By incorporating the effect of atom-atom interactions, our simulator can probe the scrambling of quantum information which is a fundamental property of black holes.
@@
Introduction:– A classical particle in the gravitational pull of a black hole cannot escape once it crosses the event horizon. However, quantum fluctuations from the interior of a black hole can tunnel across the horizon and radiate outwards, a phenomenon known as Hawking radiation [1, 2, 3, 4, 5, 6]. The emitted thermal radiation carries no information about the quantum state inside the black hole and its temperature only depends on the gravitational curvature around the horizon. The Hawking radiation can be fully analysed in the semiclassical limit ($\hbar$ small), which describes a non-interacting theory of free fields in a static background geometry [7, 8]. In addition, any quantum state entering the black hole is expected to scramble (thermalise) with the highest possible rate. This optimal ergodicity property is caused by quantum gravity interactions that are present beyond the semiclassical description of black holes, and is the subject of intense ongoing investigations [9, 10, 11, 8].
@@
Figure 1: Schematic of the black hole simulator. (a) The black hole geometry is encoded in the position-dependent tunnelling amplitudes $\kappa_{l}$ of a fermionic tight-binding chain, and change sign across the horizon. The tunnelling elements $\kappa_{l}=[f(ld)+f((l+1)d)]/(8d)$, where $d$ is the discretisation length, follow the behaviour of the black hole curvature $f(x)$, an example of which is given in (b). Insets show the local dispersion relation $E(q)$ at different positions. The left-moving branch is shifted in momentum by $\pi$ to highlight the analogy with light cones. At low energies $|E|\ll J_{0}$, $E(q)$ is linear and describes massless Dirac fermions in a curved background. (c) The light-cones collapse on each other at the horizon, $x_{h}=0$, signalling that it is impossible for classical particles to cross.
@@
There are several classical and quantum analogues that can simulate the geometric properties of a black hole, e.g. by studying sound waves propagating in a supersonic fluid [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], using ultracold atoms in the continuum [22, 23, 24, 25] or quantum circuits [26]. However, an open question remains as to how to devise a system that can produce the geometrical features of the black hole’s horizon, while also exhibiting the scrambling behaviour expected to be present in its interior. As an alternative, theoretical proposals have demonstrated how lattice Hamiltonians can simulate quantum field theories in curved space-time geometry [27, 28, 29, 30] and ultracold atoms in optical lattices have emerged as highly flexible and effective analogue quantum simulators of lattice Hamiltonians [31, 32, 33]. Indeed, optical lattices already allow us to probe intriguing quantum many-body phenomena such as entanglement entropy [34, 35], many-body localisation [36, 37, 38, 39, 40] and quantum scars [41, 42].
@@
Here, we propose to simulate black holes using a one-dimensional optical lattice and investigate the resulting quantum behaviour. We employ local Floquet driving to realise position-dependent tunnelling [43, 44] in order to encode massless Dirac fermions in the gravitational curvature of a black hole, as shown in Fig. 1. We show that wave packets of single atoms initialised inside the black hole eventually tunnel across the horizon and can be witnessed from outside the black hole as thermal Hawking radiation. Based on the effect local curvature has on the quantum evolution of the system, we propose a straightforward scheme to measure the Hawking temperature from onsite population measurements of atoms. Additionally, we demonstrate the resilience of our optical lattice simulator under a wide range of experimental imperfections. The simplicity of our scheme paves the way for various extensions. We show how our simulator is generalised to $(2+1)$D black holes by using readily available higher-dimensional optical lattices. Moreover, by introducing tunable interactions between the atoms using Feshbach resonances [45], we envisage probing the physics of black holes beyond the semiclassical regime. In this regime scrambling of information can be witnessed, thus allowing us to analyse qualitatively and quantitatively effects that are numerically or analytically very complex.
@@
Optical lattice simulator of Dirac fermions in a black hole background:– We now present the fermionic lattice Hamiltonian that in the low energy limit is mathematically equivalent to the Dirac equation in the curved space around a black hole [8, 28, 29, 46, 47, 48, 49, 50]. Our optical lattice simulator is described by a fermionic tunnelling Hamiltonian given by
$\mathcal{H}=-J_{0}\sum_{l=1}^{L}\left[\kappa_{l}(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+\hat{c}_{l+1}^{\dagger}\hat{c}_{l})\right],$ (1)
@@
where the $\hat{c}_{l}^{\dagger}$($\hat{c}_{l}$) denote the fermionic creation (annihilation) operators on site $l$, $J_{0}$ defines a global energy scale, and the $\kappa_{l}$’s are position-dependent dimensionless tunnelling amplitudes that encode the metric of the curved space. When varying the amplitudes $\kappa_{l}$ [Fig. 1(a)] according to a possible metric describing a black hole [Fig. 1(b)], the (local) dispersion relation of the lattice near zero energy mirrors the behaviour of the relativistic light-cones, see Fig. 1(c). In the continuum limit and at half filling, the behaviour of the fermionic chain can be faithfully described by the massless Dirac equation $i\gamma^{a}e^{\mu}_{a}(\partial/\partial x^{\mu}+\omega_{\mu})\psi=0$. Here the coordinates are $(x^{0},x^{1})=(ct,x)$ with $c$ the speed of light, $\gamma^{a}$ are the two-dimensional Dirac matrices, $e^{a}_{\mu}$ are the vierbeins connecting the curved metric $g_{\mu\nu}=\eta_{ab}e^{a}_{\mu}e^{b}_{\nu}$, to the flat metric $\eta_{ab}=\rm{diag}(1,-1)$, where $\omega_{\mu}$ is the spin connection and $\psi$ is the two-component spinor (see Supp. Mat. for details [51] and Refs. [52, 53] for an introduction to this formalism). The time evolution under the lattice Hamiltonian Eq. (1) is equivalent to that of a Dirac field in a curved background under the metric
@@
when $\gamma^{0}=\sigma^{z}$, $\gamma^{1}=i\sigma^{y}$, $\hbar c=J_{0}$ [54], where the spacetime curvature $f(x)$ is encoded in the lattice tunnelling amplitudes via $\kappa_{l}=[f(ld)+f((l+1)d)]/(8d)$ and $d$ is a chosen discretisation length [28, 51]. A typical Schwarzschild-like metric $f(x)$ in Eq. (2) possesses a horizon at a distance $x_{h}$ where $f(x_{h})=0$, while $f(x)<0$ describes the black hole interior ($x<x_{h}$) and $f(x)>0$ the exterior ($x>x_{h}$), as shown in Fig. 1(b). In this metric, that corresponds to the free-falling frame, the light cones compress radially near the event horizon due to the strong gravitational pull as depicted in Fig. 1(c). At the horizon, the lightcone collapses to a vertical line, signalling the impossibility for classical particles to cross. Nevertheless, quantum particles can tunnel across the horizon resulting in Hawking radiation. The change of sign of $f(x)$ across the event horizon leads to an apparent interchange of space and time coordinates within the black hole. This is encoded in the corresponding change of sign in the tunnelling amplitudes $\kappa_{l}$. A typical profile for (1+1)-dimensional black holes is $f(x)=A\tanh(x/w)$ [28, 50], where $w$ is positive and controls the steepness of the curve around the horizon and A is a scaling factor, see Fig. 1(b). For simplicity, in our proposed simulator we approximate this profile around the horizon with a linear function where the dimensionless slope $\alpha=Ad/w$ is controlled by the width $w$. This corresponds to tunnelling coefficients of the form
$\kappa_{l}=\alpha(l-l_{h}+0.5)/4,$ (3)
such that $\kappa_{l}\neq 0$ at the horizon despite the required $f(x_{h})=0$.
@@
Figure 2: A black hole simulator implemented using local Floquet driving of a 1D optical lattice. (a) The onsite potential of each site $l$ is modulated according to $\epsilon_{l}(t)=A_{l}\cos(\omega t)$ with an independent amplitude $A_{l}$ using individual tweezer beams indicated by the light-blue shading. (b) Driving amplitudes $A_{l}$ for a nine-site lattice that generate renormalised dimensionless tunnelling elements $\kappa_{l}$ following the linear profile given by Eq. (3). The inverted sign of neighbouring driving amplitudes indicates the modulations are $\pi$ out of phase. (c) The numerically calculated stroboscopic Hamiltonian $H_{S}^{t_{0}}$ for $t_{0}=0$, $\hbar\omega/\kappa J_{0}=25$, and the $A_{l}$ shown in (b).The inset shows that the stroboscopic tunnelling elements perfectly follow the target linear profile in Eq. (3) (red line). Here we took $\alpha/\kappa=0.4$ and $l_{h}=5$.
@@
The discrete lattice representation of a black hole with curvature encoded into the local tunnelling coefficients $\kappa_{l}$ in Eq. (1) can be implemented using a locally-driven, one-dimensional optical lattice, where tightly-focused far off-resonant optical tweezers [55, 56, 57, 58] independently drive the on-site potentials of individual lattice sites in an optical lattice [44], as illustrated in Fig. 2. In the following, we utilise Floquet-Bloch theory to capture the long-term evolution of the periodically driven system and demonstrate that it corresponds to a time-independent effective Hamiltonian [59, 60, 61, 62] that matches the desired position-dependent tunnelling profile given by Eq. (3). The full time-dependent Hamiltonian of the driven system is given by
$H(t)=-J_{0}\kappa\sum_{l}(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+\hat{c}_{l+1}^{\dagger}\hat{c}_{l})+\cos(\omega t)\sum_{l}A_{l}\hat{c}_{l}^{\dagger}\hat{c}_{l},$ (4)
@@
where $\kappa$ is a dimensionless value such that $\kappa J_{0}$ is the bare tunnelling between neighbouring lattice sites, $A_{l}$ is the driving amplitude of site $l$, and $\omega$ is the global driving frequency ensuring $H(t)=H(t+T)$ with period $T=2\pi/\omega$. The constant $\kappa$ is required to encode a larger range of Hawking temperatures in our simulator [51]. The $T$-periodic nature of Eq. (4) ensures that the resulting stroboscopic evolution during multiples of the global driving period $T$ can be captured by
@@
where $n\in\mathbb{N}$ and $H_{S}^{t_{0}}$ is the time-independent, “stroboscopic Hamiltonian” [61, 60]. The stroboscopic Hamiltonian is analytically difficult to calculate, but can be accessed numerically [51]. A convenient analytical approximation of $H_{S}^{t_{0}}$ can be obtained in the high-frequency regime, resulting in Hamiltonian Eq. (1) with
$\kappa_{l}=\kappa\mathcal{J}_{0}\left(\frac{|A_{l}-A_{l+1}|}{\hbar\omega}\right),$ (6)
where $\mathcal{J}_{0}(y)$ is the zeroth order Bessel function [51]. Crucially, Eq. (6) illustrates that local tunnelling elements depend only on the relative modulation between neighbouring lattice sites [44]. We can now reverse-engineer Eq. (6) to generate a sequence of $A_{l}$ and a global $\kappa$ that result in the desired tunnelling profile given in Eq. (3), and hence encode the black-hole geometry. For example, Figure 2(b) shows a series of driving amplitudes $A_{l}$ that generate $\kappa_{l}$ following the linear profile of Eq. (3) with $\alpha/\kappa=0.4$ and $l_{h}=5$. In order to minimise the absolute size of the required $A_{l}$, and thereby the resulting Floquet heating [63, 64, 61], we alternate their signs.
@@
Hawking temperature from radiation population:– To demonstrate the reliability of our simulator in reproducing the quantum features of a black hole, we first investigate the behaviour of the Hawking temperature
$T_{H}={\hbar c\over 4\pi k_{B}}\left.{df(x)\over dx}\right|_{x=x_{h}},$ (7)
@@
which depends on the curvature of the black hole at the horizon [65]. Then we propose a simple method to measure $T_{H}$ in an optical lattice experiment by utilising easily accessible in-situ density measurements to analyse the quantum walk of an initially localised atom [66, 67, 68]. In the following, we take for simplicity $\hbar=1$, $k_{B}=1$ and $c=J_{0}=1$ unless stated otherwise. This leads to $\alpha=4\pi T_{H}$ in our linearised simulator given in (3).
@@
To study the physics close to the simulated event horizon we consider the evolution of a single quantum particle, termed the walker, initially positioned inside the black hole, and its potential tunnelling across the horizon. In Fig. 3(a) we show the density evolution of such a walker under the linearised tunneling profile of Eq. (3) with $l_{h}=50$, $\alpha=1$ on a chain of $100$ sites where the walker is initially at site $l=45$, i.e. inside the black hole. The walker noticeably remains mostly confined inside the black hole but with a non-zero probability of escaping, i.e. evaporating [69, 48]. This forms the outgoing Hawking radiation. We explore walker dynamics for a range of Hawking temperatures by evolving the initially localised state |$\ket{\Psi_{i}}$ numerically using |$\ket{\Psi(t)}=e^{-i\mathcal{H}t}\ket{\Psi_{i}}$.
@@
An observer outside the horizon only has access to the visible (outside) part of the system and will therefore see the emergence of an incoherent state expected to correspond to a blackbody spectrum at temperature $T_{H}$ [65]. We compute the probability $P_{n}=\bra{E_{n}}\rho_{\textrm{out}}\ket{E_{n}}$ of a particle with energy $E_{n}$ being detected outside the horizon, where $\rho_{\textrm{out}}=\text{Tr}_{\textrm{in}}\left(\rho_{t}\right)$ is the reduced density matrix of the evolved state $\rho_{t}=\ket{\psi(t)}\bra{\psi(t)}$ | at time $t$ in the exterior of the black hole, and |$\ket{E_{n}}$ are the energy eigenstates of the Hamiltonian $H_{\rm{out}}$ taken to be acting only on the sites outside the horizon. In analogy with the particle-type solutions of the Dirac equation, we focus our analysis on positive energies, and indeed find that this spectrum decays exponentially as $P_{n}\propto e^{-E_{n}/\tilde{T}_{H}}$, as shown in Fig. 3(b). In fitting the slopes in Fig. 3(b) we ignored the exponentially suppressed contributions from very large energies as our Dirac simulation is valid in the low energy regime, and also from very small energies where discretisation effects become dominant.
@@
Figure 3: (a) Density evolution (log-scale) of a single particle initialised at $l=45$ under the Hamiltonian in Eq. (1) for a Hawking temperature $T_{H}=\alpha/(4\pi)\approx 0.08$. (b) Probability (log scale) of finding a particle with energy $E_{n}$ outside the event horizon satisfying a blackbody spectrum $P_{n}\propto e^{-E_{n}/\tilde{T}_{H}}$, evaluated at time $t=6/\alpha$, for varying $\alpha$. Extracted temperatures $\tilde{T}_{H}$ from the thermal distribution agree with the expected Hawking temperature $T_{H}$ encoded in the Hamiltonian. (c) Time evolution of the density $n_{\textrm{h}}$ at the horizon site $l_{h}=50$ for several $\alpha$ (or $T_{H}$). Same colors in (b) and (c) corresponds to the same $\alpha$. $t_{f}$ denotes the time at which peak density is reached at the horizon. (Inset) Collapsed curves with time rescaled by $1/\alpha$. (d) Hawking temperature $\tilde{T}_{H}$ extracted from spectrum in (b) against the measured time $t_{f}$ of the peak of the density at the event horizon. The solid line fits the data to the ansatz $\tilde{T}_{H}\propto t_{f}^{-1}$ (inset in log-log scale). This relationship allows for the extraction of $\tilde{T}_{H}$ from an experimentally measured peak time $t_{f}$.
@@
Experimentally, directly monitoring the populations of all eigenstates |$\ket{E_{n}}$ necessitates full state tomography in the black hole exterior, which is a tantalising task. Using the full tanh profile sketched in Fig. 1(b), it would also be possible to extract the spectrum from the experimentally accessible momentum distribution of the walker in the flat region of space. However, this would require large systems that might be challenging to realise experimentally. Hence, in the following we propose an alternative means to extract the Hawking temperature, and hence the curvature at the horizon, purely from in-situ population measurements routinely available in quantum gas microscopes [70, 71, 72, 73]. Intuitively, due to the decreasing tunnelling amplitudes close to the horizon, the walker will slow down in this region, before either being reflected back into the black hole or tunneling through to the exterior, resulting in a pronounced maximum in density that depends on $\alpha$, i.e. on the local curvature at the horizon.

### v2

@@ Abstract
In the curved spacetime of a black hole, quantum physics gives rise to distinctive effects such as Hawking radiation and maximally fast scrambling. Here, we present a scheme for an analogue quantum simulation of $(1+1)$ and $(2+1)$-dimensional black holes using ultracold atoms in a locally Floquet-driven optical lattice. We show how the effective dynamics of the driven system can generate position-dependent tunnelling amplitudes that encode the curved geometry of the black hole. Moreover, we provide a simple and robust scheme to determine the Hawking temperature of a (1+1)D simulated black hole based solely on on-site atom population measurements. Combined with the highly tunable onsite atom-atom interactions typical for cold atoms, our simulator provides a powerful and feasible platform to probe the scrambling of quantum information in black holes. We illustrate the ergodicity of our (2+1)D black-hole simulator by showing numerically that its level statistics in the hard-core limit approaches the ergodic regime faster than a globally homogeneous Hamiltonian.
I Introduction
@@
A classical particle in the gravitational pull of a black hole cannot escape once it crosses the event horizon. However, quantum fluctuations from the interior of a black hole can tunnel across the horizon and radiate outwards, a phenomenon known as Hawking radiation [1, 2, 3, 4, 5, 6]. The emitted thermal radiation carries no information about the quantum state inside the black hole, and its temperature only depends on the gravitational curvature around the horizon. This property of the black hole horizon is distinct from observed-dependent temperature phenomena, such as the Gibbon-Hawking [7, 8] and Unruh effects [9, 10]. The Hawking radiation can be fully analysed in the semiclassical limit ($\hbar$ small), which describes a non-interacting theory of free fields in a static background geometry [11, 12]. In addition, any quantum state entering the black hole is expected to scramble (thermalise) with the highest possible rate. This optimal scrambling property [13] emerges beyond the semiclassical description of black holes, and is the subject of numerous ongoing investigations [14, 15, 16, 12].
@@
Figure 1: Schematic of the black hole simulator. (a) The black hole geometry is encoded in the position-dependent tunnelling amplitudes $\kappa_{j\>j+1}$ of a fermionic tight-binding chain, and change sign across the horizon. The tunnelling elements $\kappa_{j\>j+1}=[f(jd)+f((j+1)d)]/(8d)$, where $d$ is the discretisation length, follow the behaviour of the black hole curvature $f(x)$, an example of which is given in (b). Insets show the local dispersion relation $E(q)$ at different positions. The left-moving branch is shifted in momentum by $\pi$ to highlight the analogy with light cones. At low energies $|E|\ll J_{0}$, $E(q)$ is linear and describes massless Dirac fermions in a curved background. (c) The light-cones collapse on each other at the horizon, $x_{h}=0$, signalling that it is impossible for classical particles to cross.
@@
There exist several classical and quantum analogues that can simulate the geometric properties of a black hole, e.g. by studying sound waves propagating in a supersonic fluid [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], using ultracold atoms in the continuum [28, 29, 30, 31, 32, 33], ion trapping [34] or quantum circuits [35]. However, an open question remains as to how to devise a system that can produce the geometrical features of the black hole’s horizon, while also exhibiting the scrambling behaviour expected to be present in its interior [36, 37, 38]. As an alternative, theoretical proposals have demonstrated how lattice Hamiltonians can simulate quantum field theories in curved space-time geometry [39, 40, 41, 42, 43] and ultracold atoms in optical lattices have emerged as highly flexible and effective analogue quantum simulators of lattice Hamiltonians [44, 45, 46]. Indeed, optical lattices already allow us to probe intriguing quantum many-body phenomena such as entanglement entropy [47, 48], many-body localisation [49, 50, 51, 52, 53] and quantum scars [54, 55].
@@
Here, we propose to simulate both (1+1)D and (2+1)D black holes using an optical lattice and investigate the resulting behaviour. We employ local Floquet driving to realise position-dependent tunnelling [56, 57] in order to encode massless Dirac fermions in the gravitational curvature of a black hole, as shown in Fig. 1. We show that wave packets of single atoms initialised inside the black hole in 1D eventually tunnel across the horizon and can be witnessed from outside the black hole as thermal Hawking radiation. Based on the effect local curvature has on the quantum evolution of the system, we propose a straightforward scheme to measure the Hawking temperature from onsite population measurements of atoms. Additionally, we demonstrate the resilience of our 1D optical lattice simulator under a wide range of experimental imperfections. The simplicity of our scheme paves the way for various extensions. We show how our simulator is generalised to $(2+1)$D black holes by using readily available higher-dimensional optical lattices. Introducing tuneable interactions between the atoms in an optical lattice quantum simulator can be achieved using Feshbach resonances [58]. Numerical simulations of an interacting 2D black hole across a $7\times 7$ site lattice show level statistics of a diffusive phase. Our simulator provides an opportunity to study the many-body physics of black holes beyond the semiclassical regime, where information scrambling can be witnessed.
II Optical lattice simulator of black holes
@@
We now present the fermionic lattice Hamiltonian that in the low energy limit is mathematically equivalent to the Dirac equation in the curved space around a black hole [12, 40, 41, 59, 60, 61, 62, 63]. Our optical lattice simulator is described by a fermionic tunnelling Hamiltonian given by
$H=-J_{0}\sum_{\langle i,j\rangle}\kappa_{ij}(\hat{c}_{i}^{\dagger}\hat{c}_{j}+\hat{c}_{j}^{\dagger}\hat{c}_{i}),$ (1)
@@
where the $\hat{c}_{j}^{\dagger}$($\hat{c}_{j}$) denote the fermionic creation (annihilation) operators on site $j$, $J_{0}$ defines a global energy scale, and the $\kappa_{ij}$’s are position-dependent dimensionless tunnelling amplitudes that encode the metric of the curved space. In the 1D model we present, $j\in\{1,...,N\}$ where $N$ is the number of sites. When varying the amplitudes $\kappa_{j\>j+1}$ in a 1D model [Fig. 1(a)] according to a possible metric describing a 1D black hole [Fig. 1(b)], the (local) dispersion relation of the lattice near zero energy mirrors the behaviour of the relativistic light-cones, see Fig. 1(c). In the continuum limit and at half filling, the behaviour of the fermionic chain can be faithfully described by the massless Dirac equation $i\gamma^{a}e^{\mu}_{a}(\partial_{\mu}+\omega_{\mu})\psi=0$. Here the coordinates are $(x^{0},x^{1})=(ct,x)$ with $c$ the speed of light, $\gamma^{a}$ are the two-dimensional Dirac matrices, $e^{a}_{\mu}$ are the vierbeins connecting the curved metric $g_{\mu\nu}=\eta_{ab}e^{a}_{\mu}e^{b}_{\nu}$, to the flat metric $\eta_{ab}=\rm{diag}(1,-1)$, where $\omega_{\mu}$ is the spin connection and $\psi$ is the two-component spinor (see Supp. Mat. for details [64] and Refs. [65, 66] for an introduction to this formalism). The time evolution under the lattice Hamiltonian Eq. (1) is equivalent to that of a Dirac field in a curved background under the metric
@@
when $\gamma^{0}=\sigma^{z}$, $\gamma^{1}=i\sigma^{y}$, $\hbar c=J_{0}$ [67], where the spacetime curvature $f(x)$ is encoded in the lattice tunnelling amplitudes via $\kappa_{j\>j+1}=[f(jd)+f((j+1)d)]/(8d)$ and $d$ is a chosen discretisation length [40, 64]. A typical Schwarzschild-like metric $f(x)$ in Eq. (2) possesses a horizon at a distance $x_{h}$ where $f(x_{h})=0$, while $f(x)<0$ describes the black hole interior ($x<x_{h}$) and $f(x)>0$ the exterior ($x>x_{h}$), as shown in Fig. 1(b). In this metric, that corresponds to the free-falling frame, the light cones compress radially near the event horizon due to the strong gravitational pull as depicted in Fig. 1(c). At the horizon, the lightcone collapses to a vertical line, signalling the impossibility for classical particles to cross. Nevertheless, quantum particles can tunnel across the horizon resulting in Hawking radiation. The change of sign of $f(x)$ across the event horizon leads to an apparent interchange of space and time coordinates within the black hole. This is encoded in the corresponding change of sign in the tunnelling amplitudes $\kappa_{j\>j+1}$. A typical profile for (1+1)-dimensional black holes is $f(x)=A\tanh(x/w)$ [40, 63], where $w$ is positive and controls the steepness of the curve around the horizon and A is a scaling factor, see Fig. 1(b). For simplicity, in our proposed simulator we approximate this profile around the horizon with a linear function where the dimensionless slope $\alpha=Ad/w$ is controlled by the width $w$. This corresponds to tunnelling coefficients of the form
$\kappa_{j\>j+1}=\alpha(j-j_{h}+0.5)/4,$ (3)
such that $\kappa_{j\>j+1}\neq 0$ at the horizon despite the required $f(x_{h})=0$.
@@
Figure 2: A black hole simulator implemented using local Floquet driving of a 1D optical lattice. (a) The onsite potential of each site $j$ is modulated according to $\epsilon_{j}(t)=A_{j}\cos(\omega t)$ with an independent amplitude $A_{j}$ using individual tweezer beams indicated by the light-blue shading. (b) Driving amplitudes $A_{j}$ for a nine-site lattice that generate renormalised dimensionless tunnelling elements $\kappa_{j\>j+1}$ following the linear profile given by Eq. (3). The inverted sign of neighbouring driving amplitudes indicates the modulations are $\pi$ out of phase. (c) The numerically calculated stroboscopic Hamiltonian $H_{S}^{t_{0}}$ for $t_{0}=0$, $\hbar\omega/\kappa J_{0}=25$, and the $A_{j}$ shown in (b). The inset shows that the stroboscopic tunnelling elements perfectly follow the target linear profile in Eq. (3) (red line). Here we took $\alpha/\kappa=0.4$ and $j_{h}=5$.
@@
The discrete lattice representation of a black hole with curvature encoded into the local tunnelling coefficients $\kappa_{j\>j+1}$ in Eq. (1) can be implemented using a locally-driven, one-dimensional optical lattice, where tightly-focused far off-resonant optical tweezers [68, 69, 70, 71, 72] independently drive the on-site potentials of individual lattice sites in an optical lattice [57], as illustrated in Fig. 2. In the following, we utilise Floquet-Bloch theory to capture the long-term evolution of the periodically driven system and demonstrate that it corresponds to a time-independent effective Hamiltonian [73, 74, 75, 76] that matches the desired position-dependent tunnelling profile given by Eq. (3). The full time-dependent Hamiltonian of the driven system is given by
$H(t)=-J_{0}\kappa\sum_{\langle i,j\rangle}(\hat{c}_{i}^{\dagger}\hat{c}_{j}+\hat{c}_{j}^{\dagger}\hat{c}_{i})+\cos(\omega t)\sum_{j}A_{j}\hat{c}_{j}^{\dagger}\hat{c}_{j}.$ (4)
@@
where $\kappa$ is a dimensionless value such that $\kappa J_{0}$ is the bare tunnelling between neighbouring lattice sites, $A_{j}$ is the driving amplitude of site $j$, and $\omega$ is the global driving frequency ensuring $H(t)=H(t+T)$ with period $T=2\pi/\omega$. The constant $\kappa$ is required to encode a larger range of Hawking temperatures in our simulator [64]. The $T$-periodic nature of Eq. (4) ensures that the resulting stroboscopic evolution during multiples of the global driving period $T$ can be captured by
@@
where $n\in\mathbb{N}$, and $H_{S}^{t_{0}}$ is the time-independent, “stroboscopic Hamiltonian” [75, 74]. The stroboscopic Hamiltonian is analytically difficult to calculate, but can be accessed numerically [64]. A convenient analytical approximation of $H_{S}^{t_{0}}$ can be obtained in the high-frequency regime, resulting in Hamiltonian Eq. (1) with
$\kappa_{ij}=\kappa\mathcal{J}_{0}\left(\frac{|A_{i}-A_{j}|}{\hbar\omega}\right),$ (6)
where $\mathcal{J}_{0}(y)$ is the zeroth order Bessel function [64]. Crucially, Eq. (6) illustrates that local tunnelling elements depend only on the relative modulation between neighbouring lattice sites [57]. We can now reverse-engineer Eq. (6) to generate a sequence of $A_{j}$ and a global $\kappa$ that result in the desired tunnelling profile given in Eq. (3), and hence encode the black-hole geometry. For example, Figure 2(b) shows a series of driving amplitudes $A_{j}$ that generate $\kappa_{j\>j+1}$ following the linear profile of Eq. (3) with $\alpha/\kappa=0.4$ and $j_{h}=5$. In order to minimise the absolute size of the required $A_{l}$, and thereby the resulting Floquet heating [77, 78, 75], we alternate their signs.
@@ II.1 Hawking temperature from radiation population
To demonstrate the reliability of our simulator in reproducing the quantum features of a black hole, we first investigate the behaviour of the Hawking temperature
$T_{H}=\frac{\hbar c}{4\pi k_{B}}\left.\frac{df(x)}{dx}\right|_{x=x_{h}},$ (7)
@@
which depends on the curvature of the black hole at the horizon [3]. We then propose a simple method to measure $T_{H}$ in an optical lattice experiment by utilising easily accessible in-situ density measurements to analyse the quantum walk of an initially localised atom [79, 80, 81]. In the following, we take for simplicity $\hbar=1$, $k_{B}=1$ and $c=J_{0}=1$ unless stated otherwise. This leads to $\alpha=4\pi T_{H}$ in our linearised simulator given in (3).
@@
To study the physics close to the simulated event horizon we consider the evolution of a single quantum particle, termed the walker, initially positioned inside the black hole, and its potential tunnelling across the horizon. In Fig. 3(a) we show the density evolution of such a walker under the linearised tunnelling profile of Eq. (3) with $j_{h}=50$, $\alpha=1$ on a chain of $100$ sites where the walker is initially at site $j=45$, i.e. inside the black hole. The walker noticeably remains mostly confined inside the black hole but with a non-zero probability of escaping, i.e. evaporating [82, 61]. This forms the outgoing Hawking radiation. We explore walker dynamics for a range of Hawking temperatures by evolving the initially localised state |$\ket{\Psi_{i}}$ numerically using |$\ket{\Psi(t)}=e^{-i\mathcal{H}t}\ket{\Psi_{i}}$.
@@
An observer outside the horizon only has access to the visible (outside) part of the system and will therefore see the emergence of an incoherent state expected to correspond to a blackbody spectrum at temperature $T_{H}$ [7]. We compute the probability $P_{n}=\bra{E_{n}}\rho_{\textrm{out}}\ket{E_{n}}$ of a particle with energy $E_{n}$ being detected outside the horizon, where $\rho_{\textrm{out}}$ is the density matrix of the evolved state $\rho_{t}=\ket{\psi(t)}\bra{\psi(t)}$ | at time $t$ truncated at the event horizon, and |$\ket{E_{n}}$ are the energy eigenstates of the Hamiltonian $H_{\rm{out}}$ taken to be acting only on the sites outside the horizon. In analogy with the particle-type solutions of the Dirac equation, we focus our analysis on positive energies, and indeed find that this spectrum decays exponentially as $P_{n}\propto e^{-E_{n}/\tilde{T}_{H}}$, as shown in Fig. 3(b).
In fitting the slopes in Fig. 3(b) we ignored the exponentially suppressed contributions from very large energies as our Dirac simulation is valid in the low energy regime. We also implemented a cut-off of contributions from very small energies where discretisation effects become dominant. While large wavevectors can lead to greybody effects distorting the blackbody spectrum, the thermal signature remains robust for the energy scale chosen in Fig. 3(b) [83].
@@
Figure 3: (a) Density evolution (log-scale) of a single particle initialised at $j=45$ in a chain of length $L=100$ under the Hamiltonian in Eq. (1) using a Hawking temperature $T_{H}=\alpha/(4\pi)\approx 0.08$. (b) Probability (log scale) of finding a particle with energy $E_{n}$ outside the event horizon satisfying a blackbody spectrum $P_{n}\propto e^{-E_{n}/\tilde{T}_{H}}$, evaluated at time $t=6/\alpha$, for varying $\alpha$. Extracted temperatures $\tilde{T}_{H}$ from the thermal distribution agree with the expected Hawking temperature $T_{H}$ encoded in the Hamiltonian. (c) Time evolution of the density $n_{\textrm{h}}$ at the horizon site $j_{h}=50$ for several $\alpha$ (or $T_{H}$). Same colors in (b) and (c) corresponds to the same $\alpha$. $t_{f}$ denotes the time at which peak density is reached at the horizon. (Inset) Collapsed curves with time rescaled by $1/\alpha$. (d) Hawking temperature $\tilde{T}_{H}$ extracted from spectrum in (b) against the measured time $t_{f}$ of the peak of the density at the event horizon. The solid line fits the data to the ansatz $\tilde{T}_{H}\propto t_{f}^{-1}$ (inset in log-log scale). This relationship allows for the extraction of $\tilde{T}_{H}$ from an experimentally measured peak time $t_{f}$.
@@
Experimentally, directly monitoring the populations of all eigenstates |$\ket{E_{n}}$ necessitates full state tomography in the black hole exterior, which is a tantalising task. Using the full tanh profile sketched in Fig. 1(b), it would also be possible to extract the spectrum from the experimentally accessible momentum distribution of the walker in the flat region of space. However, this would require large systems that might be challenging to realise experimentally. Hence, in the following we propose an alternative means to extract the Hawking temperature, and hence the curvature at the horizon, purely from in-situ population measurements routinely available in quantum gas microscopes [84, 85, 86, 87]. Intuitively, due to the decreasing tunnelling amplitudes close to the horizon, the walker will slow down in this region, before either being reflected back into the black hole or tunnelling through to the exterior, resulting in a pronounced maximum in density that depends on $\alpha$, i.e. on the local curvature at the horizon.